OpenStudy (shamil98):

What is the sum of the two vectors (-1,-4) and (3,5)?

OpenStudy (amistre64):

all you do is add up like components ...... given (a,b) and (p,q) (a,b) + (p,q) --------- (a+p, b+q)

OpenStudy (shamil98):

So... it would be -1 + 3 = 2 -4 + 5 = 1 so (2,1)

OpenStudy (amistre64):

looks about right, i perfer a vertical setup meself (-1,-4) +( 3, 5) --------- ( 2 , 1 )